# Strict comparison for twisted group $\mathrm{C}^{*}$-algebras

### v1

@@ 1. Introduction
More recently, strict comparison has attracted significant interest in the context of $C^{*}$-algebras arising from combinatorial or algebraic data. This includes not only the simple, nuclear case —where comparison properties are tightly connected to classification results— but also the simple, non-nuclear setting [AGKEP24, KES25, Vig25], as well as certain non-simple but nuclear algebras [EV25, HRW07, RT17].
@@
This paper focuses on reduced group $C^{*}$-algebras $C^{*}_{\mathrm{red}}(G)$ and their twisted counterparts $C^{*}_{\mathrm{red}}(G,\sigma)$ for non-amenable groups $G$. In the untwisted case, a recent breakthrough in [AGKEP24] shows that a broad class of simple group $C^{*}$-algebras are selfless, and therefore pure and in particular possess strict comparison. Specifically, these results apply to reduced group $C^{*}$-algebras of selfless groups with the rapid decay property —which include all finitely generated, acylindrically hyperbolic groups with trivial finite radical and rapid decay by [AGKEP24, Theorem 3.3]. See also [Vig25], which proves that cocompact lattices in $\mathrm{PSL}(n,K)$ for $K$ a local field of characteristic zero are selfless.
@@
Let $G$ be a finitely generated, acylindrically hyperbolic group with rapid decay, and let $\sigma\in Z^{2}(G,\mathbb{T})$. Then $C^{*}_{\mathrm{red}}(G,\sigma)$ is pure. In particular, it has strict comparison.
It is worth noting that Theorem B covers all nonelementary, hyperbolic groups, also allowing for a nontrivial finite radical; see Corollary 2.3. As a concrete example, we obtain that the reduced group $C^{*}$-algebra $C^{*}_{\mathrm{red}}(\operatorname{GL}(2,{\mathbb{Z}}))$ is pure. Indeed, the group $\operatorname{GL}(2,\mathbb{Z})$ is virtually free and therefore hyperbolic, but its centre witnesses a nontrivial finite radical and so it was not covered by the results of [AGKEP24].
@@ Acknowledgments
This work was carried out when SR visited HT and EV at Chalmers University of Technology. They are grateful to the institution for its hospitality. EV and HT also wish to thank David Jekel for discussions on the subject.
@@ 2. The proofs
Let $G$ be a finitely generated, acylindrically hyperbolic group with rapid decay, and let $\sigma\in Z^{2}(G,\mathbb{T})$. We need to show that $C^{*}_{\mathrm{red}}(G,\sigma)$ is pure.
@@
It follows that $C^{*}_{\mathrm{red}}(G,\sigma)$ is a finite direct sum of matrix amplifications of pure $C^{*}$-algebras, and hence pure itself. ∎
@@
By definition, every hyperbolic group is finitely generated. Further, as noted in [Osi16, Appendix 8], nonelementary, hyperbolic groups are acylindrically hyperbolic. Further, hyperbolic groups enjoy the rapid decay property by [Jol90, dlH88]. Thus, Theorem B applies. ∎

### v2

@@ 1. Introduction
More recently, strict comparison has attracted significant interest in the context of $C^{*}$-algebras arising from combinatorial or algebraic data. This includes not only the simple, nuclear case —where comparison properties are tightly connected to classification results— but also the simple, non-nuclear setting [AGKEP24, KES25, Vig25, HKER25], as well as certain non-simple but nuclear algebras [EV25, HRW07, RT17].
@@
This paper focuses on reduced group $C^{*}$-algebras $C^{*}_{\mathrm{red}}(G)$ and their twisted counterparts $C^{*}_{\mathrm{red}}(G,\sigma)$ for non-amenable groups $G$. In the untwisted case, a recent breakthrough in [AGKEP24] shows that a broad class of simple group $C^{*}$-algebras are selfless, and therefore pure and in particular possess strict comparison. Specifically, these results apply to reduced group $C^{*}$-algebras of selfless groups with the rapid decay property —which include all finitely generated, acylindrically hyperbolic groups with trivial finite radical and rapid decay by [AGKEP24, Theorem 3.3]. See also [Vig25], which proves that cocompact lattices in $\mathrm{PSL}(n,K)$ for $K$ a local field are selfless, which in characteristic zero combines with results on the rapid decay property [RRS98, Laf00].
@@
Let $G$ be a countable, acylindrically hyperbolic group with rapid decay, and let $\sigma\in Z^{2}(G,\mathbb{T})$. Then $C^{*}_{\mathrm{red}}(G,\sigma)$ is pure. In particular, it has strict comparison.
It is worth noting that Theorem B covers all nonelementary, hyperbolic groups, also allowing for a nontrivial finite radical; see Corollary 2.3. As a concrete example, we obtain that the reduced group $C^{*}$-algebra $C^{*}_{\mathrm{red}}(\operatorname{GL}(2,{\mathbb{Z}}))$ is pure. Indeed, the group $\operatorname{GL}(2,\mathbb{Z})$ is virtually free and therefore hyperbolic, but its centre witnesses a nontrivial finite radical and so it was not covered by the results of [AGKEP24]. We discuss further examples of pure twisted group C*-algebras in Example 2.4 and Example 2.5.
@@ Acknowledgments
This work was carried out when SR visited HT and EV at Chalmers University of Technology. They are grateful to the institution for its hospitality. EV and HT also wish to thank David Jekel for discussions on the subject. The authors thank Itamar Vigdorovich for helpful comments on a first version of this work.
@@ 2. The proofs
Assume first that $G$ is finitely generated, that is, let $G$ be a finitely generated, acylindrically hyperbolic group with rapid decay, and let $\sigma\in Z^{2}(G,\mathbb{T})$. We show that $C^{*}_{\mathrm{red}}(G,\sigma)$ is pure.
@@
It follows that $C^{*}_{\mathrm{red}}(G,\sigma)$ is a finite direct sum of matrix amplifications of pure $C^{*}$-algebras, and hence pure itself.
Now let $G$ be any countable, acylindrically hyperbolic group with rapid decay. By [Osi16, Theorem 1.2], there is a non-elementary acylindrical action of $G$ on a hyperbolic space $X$ and $G$ is not virtually cyclic. So by [Osi16, Theorem 1.1], there are infinitely many pairwise independent loxodromic elements in $G$. Fix two such elements $g,h\in G$. Given an arbitrary finite subset $F\subseteq G$, the group $H=\langle S\cup\{g,h\}\rangle$ acts acylindrically hyperbolic on $X$ and its limit set contains the limit set of $\langle g,h\rangle$. So the action of $H$ on $X$ is non-elementary. This shows that $H$ is acylindrically hyperbolic. Since rapid decay passes to arbitrary subgroups by [Jol90, Proposition 2.1.1], we have shown that $G$ is the directed union of finitely generated, acylindrically hyperbolic subgroups with rapid decay. Using that pureness passes to inductive limits [PTV25, Theorem D], the result follows from the finitely generated case. ∎
@@
By definition, every hyperbolic group is finitely generated. As noted in [Osi16, Appendix 8], nonelementary, hyperbolic groups are also acylindrically hyperbolic. Further, hyperbolic groups enjoy the rapid decay property by [Jol90, dlH88]. Thus, Theorem B applies. ∎
While free groups do not have any non-trivial twisted group C*-algebras due to the fact that $\mathrm{H}^{2}(\mathbb{F}_{n},\mathbb{T})$ vanishes, many other groups to which Theorem A applies have non-trivial second cohomology with coefficients in the circle group.
Example 2.4.
Denote by $\Sigma_{g}$ the closed surface of genus $g\geq 2$ and $G=\pi_{1}(\Sigma_{g})$ the associated surface group. Then $\Sigma_{g}$ is torsion-free hyperbolic. Its homology groups are $\mathrm{H}_{0}(G,{\mathbb{Z}})\cong{\mathbb{Z}}$, $\mathrm{H}_{1}(G,{\mathbb{Z}})\cong{\mathbb{Z}}^{2g}$ and $\mathrm{H}_{2}(G,{\mathbb{Z}})\cong{\mathbb{Z}}$. So the universal coefficient theorem yields an isomorphism $\mathrm{H}^{2}(G,\mathbb{T})\cong\mathrm{Hom}(\mathbb{Z},\mathbb{T})\cong\mathbb{T}$. For $\theta\in\mathbb{T}$, we denote by $\sigma_{\theta}\in\mathrm{H}^{2}(G,\mathbb{T})$ the unique 2-cocycle mapping to $\theta$. Then $C^{*}_{\mathrm{red}}(G,\sigma_{\theta})$ is selfless by Theorem A.
The next example shows how Vigdorovich’s work on cocompact lattices in projective special linear groups [Vig25] can be lift to special linear groups. We remark that cocompact lattices in $\mathrm{SL}(3,K)$ can have a non-trivial centre, so that $C^{*}_{\mathrm{red}}(G)$ cannot be simple, hence not selfless.
Example 2.5.
Let $K$ be a local field of characteristic zero and $G\leq\mathrm{SL}(3,K)$ a cocompact lattice. Given any $\sigma\in\mathrm{H}^{2}(G,\mathbb{T})$, we show that $C^{*}_{\mathrm{red}}(G,\sigma)$ is pure. Denote by $\mathrm{Ad}\colon\mathrm{SL}(3,K)\longrightarrow\mathrm{PSL}(3,K)$ the adjoint representation. The kernel of $\mathrm{Ad}$ equals the centre of $\mathrm{SL}(3,K)$. It has order at most $3$, since there are at most three third roots of unity in $K$. Write $Z=\ker(\mathrm{Ad}|_{G})$. Then $Z$ is a finite, central subgroup of $G$. The quotient $\mathrm{Ad}(G)\cong G/Z$ and all its finite index subgroups are cocompact lattices in $\mathrm{PSL}(3,K)$. So by [Vig25, Theorem 1.2 and Lemma 2.3], they are selfless. Further, they have the rapid decay property by [RRS98, Laf00]. We can now argue as in the proof of Theorem B.